# Collaborative Agents for Software Engineering

### v2

@@ Abstract
Code review is a heavily collaborative process, which aims at ensuring the overall quality and reliability of software. While it provides massive benefits, the implementation of code review in an organization faces several challenges that make its automation appealing. Automated code review tools have been around for a while and are now improving thanks to the adoption of novel AI models, which help can learn about standard practices and systematically check that the reviewed code adheres to them. Unfortunately, existing methods fall short: they often target a single input-output generative model, which cannot simulate the collaboration interactions in code review to account for various perspectives; they are also sub-performing on various critical code review sub-tasks. In this paper, we advance the state of the art in code review automation by introducing CodeAgent, a novel multi-agent-based system for code review. Fundamentally, CodeAgent is steered by QA-Checker (short for “Question-Answer Checking"), a supervision agent, designed specifically to ensure that all agents’ contributions remain relevant to the initial review question. CodeAgent is autonomous, multi-agent, and Large language model-driven. To demonstrate the effectiveness of CodeAgent, we performed experiments to assess its capabilities in various tasks including 1) detection of inconsistencies between code changes and commit messages, 2) detection of vulnerability introduction by commits, and 3) validation of adherence to code style. Our website is accessed in https://code-agent-new.vercel.app/index.html.
@@ 1 Introduction
Code review, as a fundamental activity in software engineering, has been widely studied in the literature [2; 3; 8]. It involves multiple team members, with different expertise and experience, collaborating to check the code along several dimensions, including whether the new code is aligned with existing code in terms of style [14], whether the pull request contents are consistent [43], whether some vulnerability is being injected [4], etc. This review effort is critical for ensuring the stability, quality and readability of the code. Historically, code review has been an intensive manual exercise. Nowadays, however, software projects evolve at a rapid pace with a high frequency of commits and pull requests. For example, in 2022, GitHub, a global platform for software development, witnessed over 3.5 billion contributions, including commits, pull requests, and other forms of contributions [16; 30]. Reviewing these contributions before they are integrated into any code base requires automated tools to ease the workload of code contributors and code reviewers.
In recent literature, various approaches [35; 32; 34] have been proposed to elevate code review automation, using Transformers [36] for code modeling, code abstraction tools such as src2abs [33] and code transformation engines such as AutoTransform [32]. These state-of-the-art approaches however mainly focused on how to rewrite and adapt the submitted code. Code review, however, is actually a collaborative and iterative process involving multiple stakeholders exchanging on a specific thought process [2] and addressing a wide variety of subtasks such as analyzing the consistency between a code change and the associated commit message () [43], detecting potential vulnerability injection () [4], assessing the consistency of code format () [14], etc.
To address the challenges of complex software collaboration eco-system, researchers have applied agent-based systems across a spectrum of tasks, advancing past the conventional single input-output paradigm [40; 38]. Notably, the advent of multi-agent collaboration stands out as a key innovation, especially in simulating human-like behaviors [10; 24; 27] and harnessing the collective strengths of various agents [7; 22; 17].
@@
In this paper, drawing on the success of agent-based collaboration, we developed an agent-based framework, CodeAgent, to simulate the dynamics of a collaborative team engaged in code review processes, incorporating various roles such as code authors, reviewers, and decision-makers. Nonetheless, a notable hurdle within multi-agent systems and Chain-of-Thought (CoT) reasoning is the propensity for conversation topics to drift off course, underlining the necessity for strategies to maintain topic relevance and coherence [20; 6]. The occurrence of drift, often triggered by the model-inspired tangents or the randomness of Large Language Models (LLMs), necessitating the integration of a QA-Checker. This QA-Checker, serving as a crucial supervisory agent within CodeAgent, meticulously monitors the flow of conversation, ensuring that questions and responses remain pertinent and on track, thus maintaining the dialogue’s intended direction [20; 6]. As an instruction-driven entity, the QA-Checker not only refines queries but also realigns answers to match the original intent, employing a systematic approach grounded in a mathematical framework. This involves leveraging a quality assessment function and the Newton-Raphson optimization method [41] to iteratively guide the conversation towards optimal coherence and relevance.
To evaluate the performance of CodeAgent, we employ two distinct sources of data. The first encompasses pre-existing datasets: Trans-Review, AutoTransform, and T5-Review, which have been referenced by leading research in the field. By conducting experiments with these datasets, we aim to compare the code revision capabilities of CodeAgent against those identified as state-of-the-art. The second data source comprises a collection of Pull Requests from GitHub, featuring an extensive array of commits, messages, and comments across nine programming languages post-April 2023. These languages include Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. This diverse dataset is specifically curated for testing format analysis. The experimental results indicate that CodeAgent outperforms the state-of-the-art results in terms of qualitative metrics for code review, achieving a 41.54 percentage point (pp) increase in hit rate (Rate in Table 3) confirming vulnerable issues within merged commits across the nine programming languages. CodeAgent also excels in consistency detection, vulnerability identification, and format analysis, outperforming ChatGPT in overall metrics.
In conclusion, we summarize our contributions as follows:
We build a code review dataset with more than 3,545 real-world commits, commit messages, and corresponding original files for assessing code review model performance in various areas including consistency detection between commits and commit messages, vulnerability detection, code style detection, and code revision.
@@
Experimental results show that CodeAgent confirms 104 (23.20% higher) more vulnerable issues. Compared to state-of-the-art, CodeAgent improves 5.62pp and 4.00pp of recall and F1-Score, respectively, for the consistency detection between commit and commit message. CodeAgent also boosts 15.96pp and 10.45pp of the recall score and F1-Score, respectively, for the format consistency detection. On the code revision task, CodeAgent surpasses the state-of-the-art from 29.80pp to 31.60pp of the Edit Progress (EP) metric [44] on average.
2 Tasks and Definition
In this section, we summarize our tasks (❶,❷,❸,and ❹) and definitions (,,) in Table. 1.
| Index | Tasks |
| ❶ | Semantic consistency detection between commit and commit message |
| ❷ | Vulnerability analysis |
| ❸ | Format consistency detection |
| ❹ | Code revision |
| Short Term | Definition |
| Consistency analysis between commit and commit message | |
| Vulnerability analysis | |
| Format consistency analysis between commit and original files |
@@ 3 CodeAgent
This section details the methodology behind our innovative CodeAgent framework. We first describe the defined role cards in Section 3.1 and discuss the pipeline in Section 3.2. Finally, we will discuss the design of the QA-Checker.
@@ 3.1 Role Card Definition
As shown in Figure 1, we defined six characters in our simulation system (CodeAgent), including , , , , , , and they are defined for different specific tasks.
For each, we defined a role card, which contains: 1) The role name is put on the left-upper corner of each card; 2) The phases of the role involved are put on the right-upper corner of each card; 3) On each role card, we show the role-involved conversation and collaborative roles; 4) We show the intermediate output of the role on the right-hand side of the card; and 5) Finally, we put the corresponding files or content out of conversations on the bottom of the card.
All tasks are processed by the collaborative work of two agents in their multi-round conversations. For example, as a role , her responsibility is to do the code review for given codes and files in three aspects (task ❶, ❷, and ❸ in Table 1) and provide a detailed description of observation. ’s code review activity is under the assistance with as shown in Figure 2. Meanwhile, with the Review’s assistance, can process the code revision as shown in the ‘Revised codes’ part in the card in Figure 1. Apart from , also cooperates with and in the simulated team.
3.2 Pipeline
We specifically consider the following sub-tasks of code review: consistency detection between commit and commit message () [43], vulnerability injection detection () [4], Format consistency detection () [14], and code revision [44]. We defined six characters and four phases for the framework. The roles of characters are demonstrated in Figure 1. Each phase contains multiple conversations and each conversation happens between agents. The four phases consist of ① Basic Info Sync, containing the roles of chief executive officer (), chief technology officer () and Coder to conduct modality and language analysis; ② Code Review, leveraging the and for actual code review (i.e., target sub-tasks); ③ Code Alignment, supporting the and to correct the commit by code revision or suggestion feedback to the author; ④ Document, finalizing by synthesizing the opinions of the , , and to give out the final comments.
Apart from six defined roles, the proposed architecture of CodeAgent consists of phase-level and conversation-level components. At the phase level, the waterfall model is used to break down the code review process into four sequential phases. At the conversation level, each phase is further divided into atomic conversations. These atomic conversations involve task-oriented role-playing between two agents, promoting collaborative communication. One agent works as the instructor and the other works as the assistant. Communication follows an instruction-following style, where agents interact to accomplish a specific subtask within each conversation and each conversation is under the supervision of QA-Checker. QA-Checker is used to align the consistency of questions and answers between the instructor and assistant in a conversation to avoid digression. QA-Checker will be introduced in Section 3.3.
Here, we take an example to show the pipeline of CodeAgent. As shown in Figure 2, CodeAgent receives the request to do the code review with the submitted commit, commit message, and original files. In the first phase, , , and will cooperate to recognize the modality of the input (e.g., document, code) and language (e.g., Python, Java, and Go). In the second phase, with the help of , will write an analysis report about consistency analysis, vulnerability analysis, format analysis, and suggestions for code revision. Then, in the third phase, according to analysis reports, will align or revise the code if it finds incorrect snippets with the help of . cooperates with and to summarize the document and codes about the whole code review in the final phase.
@@ 3.3 Self-Improving CoT with QA Checker
QA-Checker is an instruct-driven agent, designed to fine-tune the question inside a conversation to drive the generated answer related to the question. As shown in Figure 3, the initial question (task instruction) is represented as , and the first answer of the conversation between and is represented as . If QA-Checker identifies that is inappropriate for , it generates additional instructions attached to the original question (task instruction) and combines them to ask agents to further generate a different answer. The combination in Figure 3 is defined as , where is the additional instruction attached. The conversation between two agents is held until the generated answer is judged as appropriate by QA-Checker, it reaches the maximum dialogue times, otherwise.
@@
The QA-Checker is an instruction-driven agent, crucial in refining questions and answers within a conversation to ensure relevance and precision. Its operation can be understood through the following lemma and proof.
Lemma 3.1.
Let denote the quality assessment function of the QA-Checker for question-answer pair in a conversation at the -th iteration. Assume is twice differentiable and its Hessian matrix is positive definite. If the QA-Checker modifies the question to by attaching an additional instruction , and this leads to a refined answer , then the sequence converges to an optimal question-answer pair , under specific regularity conditions.
Proof.
The QA-Checker refines the question and answers using the rule:
Substituting the update rule and rearranging, we get:
In practical terms, this lemma and its proof underpin the QA-Checker’s ability to refine answers iteratively. The QA-Checker assesses the quality of each answer concerning the posed question, employing advanced optimization techniques that is modeled by the modified Newton-Raphson method to enhance answer quality. This framework ensures that, with each iteration, the system moves closer to the optimal answer, leveraging both first and second-order derivatives for efficient and effective learning.
4 Experimental Design
We evaluate the performance of CodeAgent in various qualitative and quantitative experiments in nine programming languages, on four metrics. In this Section, we will discuss experimental settings, including datasets, metrics, and baselines. For more information, please see Appendix B.
@@ 4.1 Datasets
For our research, we leverage datasets from prior studies as referenced in the state-of-the-art [44]. Specifically, we utilized Trans-Review, AutoTransform, and T5-Review to assess the edit progress (EP) metric which is also designed by them [44].
As shown in Table 2, in terms of new data collection, which is called codeData, is collected using the GitHub REST API, encompasses over 3,545 commits and 2,933 pull requests from more than 180 projects in nine programming languages (Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby). It focuses on consistency and format detection, featuring both positive and negative samples segmented by the merged and closed status of pull requests across various languages. The detailed information about the dataset can be seen in Appendix-Section E.
@@ 4.2 Metrics
F1-Score and Recall. We utilized the F1-Score and recall to evaluate our method’s effectiveness. The F1-Score, a balance between precision and recall, is crucial for distinguishing between false positives and negatives. Recall measures the proportion of actual positives correctly identified [19].
Edit Progress (EP). EP evaluates the improvement in code transitioning from erroneous to correct by measuring the reduction in edit distance between the original code and the prediction. A higher EP indicates better efficiency in code generation [9; 11; 44].
Hit Rate (Rate) We also use hit rate to evaluate the rate of confirmed vulnerable issues out of the found issues by approaches.
@@ 4.3 State-of-the-Art Tools and Models
Our study evaluates various tools and models for code revision and modeling. Trans-Review [35] employs src2abs for code abstraction, effectively reducing vocabulary size. AutoTransform [32] uses Byte-Pair Encoding for efficient vocabulary management in pre-review code revisions. T5-Review [34] leverages the T5 architecture, emphasizing improvement in code review through pre-training on code and text data. In handling both natural and programming languages, CodeBERT [12] adopts a bimodal approach, while GraphCodeBERT [13] incorporates code structure into its modeling. CodeT5 [37], based on the T5 framework, is optimized for identifier type awareness, aiding in generation-based tasks. Additionally, we compare these tools with ChatGPT [26] by OpenAI, notable for its human-like text generation capabilities in natural language processing.

### v4

# CodeAgent: Autonomous Communicative Agents for Code Review

@@ Abstract
Code review, which aims at ensuring the overall quality and reliability of software, is a cornerstone of software development. Unfortunately, while crucial, Code review is a labor-intensive process that the research community is looking to automate. Existing automated methods rely on single input-output generative models and thus generally struggle to emulate the collaborative nature of code review. This work introduces CodeAgent, a novel multi-agent Large Language Model (LLM) system for code review automation. CodeAgent incorporates a supervisory agent, QA-Checker, to ensure that all the agents’ contributions address the initial review question. We evaluated CodeAgent on critical code review tasks: (1) detect inconsistencies between code changes and commit messages, (2) identify vulnerability introductions, (3) validate code style adherence, and (4) suggest code revisions. The results demonstrate CodeAgent’s effectiveness, contributing to a new state-of-the-art in code review automation. Our data and code are publicly available (https://github.com/Code4Agent/codeagent).
[…]
CodeAgent: Autonomous Communicative Agents for Code Review
@@ 1 Introduction
Code review Bacchelli and Bird (2013); Bosu and Carver (2013); Davila and Nunes (2021) implements a process wherein software maintainers examine and assess code contributions to ensure quality and adherence to coding standards, and identify potential bugs or improvements. In recent literature, various approaches Tufano et al. (2021, 2022) have been proposed to enhance the performance of code review automation. Unfortunately, major approaches in the field ignore a fundamental aspect: the code review process is inherently interactive and collaborative Bacchelli and Bird (2013). Instead, they primarily focus on rewriting and adapting the submitted code Watson et al. (2022); Thongtanunam et al. (2022); Staron et al. (2020). In this respect, an effective approach should not only address how to review the submitted code for some specific needs (e.g., vulnerability detection). Still, other non-negligible aspects of code review should also be considered, like detecting issues in code formatting or inconsistencies in code revisions. However, processing multiple sub-tasks requires interactions among different staff members in a real code review scenario, which makes it chellenging to design a bot that performs code review automatically.
Agent-based systems are a newly emerging paradigm and a computational framework in which autonomous entities (aka agents) interact with each other Li et al. (2023a); Qian et al. (2023); Hong et al. (2023) to perform a task. Agent-based approaches have been proposed to address a spectrum of software engineering tasks Qian et al. (2023); Zhang et al. (2024), moving beyond the conventional single input-output paradigm due to their remarkable ability to simulate and model complex interactions and behaviors within dynamic environments Xi et al. (2023); Wang et al. (2023). Recently, with multi-agent systems, the strengths of various agents are utilized to simulate human-like decision-making processes Du et al. (2023); Liang et al. (2023); Park et al. (2023), improving performance in many tasks Chen et al. (2023); Li et al. (2023b); Hong et al. (2023). This paradigm fits well with the challenge of code review, which involves multiple reviewers interacting based on their diverse skills and roles to achieve a thorough review of code.
@@
This paper. Drawing from the success of agent-based collaboration, we propose a multi-agent-based framework CodeAgent to simulate the dynamics of a collaborative team engaged in the code review process, incorporating various roles such as code change authors, reviewers, and decision makers. In particular, as a major contribution towards CodeAgent, we address the challenge of prompt drifting Zheng et al. (2024); Yang et al. (2024), commonly found in multi-agent systems and Chain-of-Thought (CoT) reasoning. This issue, characterized by conversations that stray from the main topic, highlights the need for strategies to keep discussions on track and coherent Greyling (2023); Chae et al. (2023). This drift, often triggered by the model-inspired tangents or the randomness of Large Language Models (LLMs), necessitates the integration of a supervisory agent. We employ an agent named QA-Checker (for "Question-Answer Checker") that monitors the flow of conversation, ensuring that questions and responses remain relevant and on track, thus maintaining the dialogue’s intended objective. Such an agent not only refines queries but also realigns answers to match the original intent, employing a systematic approach grounded in a mathematical framework.
To evaluate the performance of CodeAgent, we first assess its effectiveness for typical review objectives such as finding vulnerabilities 4.1 and validating the consistency and alignment of the code format 4.2. We also compare CodeAgent with state-of-the-art generic and code-specific language models Feng et al. (2020); OPENAI (2022). Finally, we assess the performance of CodeAgent compared to the state of the art for code revision suggestions Tufano et al. (2021); Thongtanunam et al. (2022); Tufano et al. (2022). Since each of these related works presents a specific dataset, we also employ them toward a fair comparison. At the same time, we also collect pull requests from GitHub, featuring an extensive array of commits, messages, and comments to evaluate advanced capabilities.The experimental results display that CodeAgent significantly outperforms the state-of-the-art, achieving a 41% increase in hit rate for detecting vulnerabilities. CodeAgent also excels in consistency checking and format alignment, outperforming the target models. Finally, CodeAgent showcases its robustness for code revision by demonstrating superior average edit progress.
We summarize our contributions as follows:
@@
We build a new dataset with about than 3 500 real-world code changes and commit messages. This dataset, which includes all relevant files and details in a standalone manner, is useful for assessing advanced code review tasks such as vulnerability detection, code style detection, and code revision suggestions.
We demonstrate the effectiveness of the QA-Checker. This agent monitors the conversation flow to ensure alignment with the original intent, addressing the common prompt drifting issues in multi-agent systems.
Experimental evaluation highlights the performance of CodeAgent: In terms of vulnerability detection, CodeAgent outperforms GPT-4 and CodeBERT by 3 to 7 percentage points in terms of the number of vulnerabilities detected. In terms of format alignment, CodeAgent outperforms ReAct by about 14% in terms of recall of inconsistency detection. On the code revision task, CodeAgent surpasses the state of the art in the software engineering literature: in terms of the Edit Progress metric Zhou et al. (2023), CodeAgent offers an average performance improvement of about 30%.
@@ 2 CodeAgent
This section details the methodology behind our novel CodeAgent framework. We discuss tasks and definition in Sec 2.1, pipeline in Section 2.2, defined role cards in Section 2.3, and the design of the QA-Checker in Sec 2.4.
2.1 Tasks
We define CA, VA, FA, and CR in as following:
CA Zhang et al. (2022): Consistency analysis between code change and commit message; the task is to detect cases where the commit message accurate describes (in natural language) the intent of code changes (in programming language).
VA Braz et al. (2022): Vulnerability analysis; the task is to identify cases where the code change introduces a vulnerability in the code.
FA Han et al. (2020): Format consistency analysis between commit and original files; the task is to validate that the code change formatting style is not aligned with the target code.
CR Zhou et al. (2023): Code revision; this task attempts to automatically suggest rewrites of the code change to address any issue discovered.
2.2 Pipeline
We defined six characters and four phases for the framework. The roles of the characters are illustrated in Figure 1. Each phase contains multiple conversations, and each conversation happens between agents. The four phases consist of 1) Basic Info Sync, containing the roles of chief executive officer (CEO), chief technology officer (CTO), and Coder to conduct modality and language analysis; 2) Code Review, leveraging the Coder and Reviewer for actual code review (i.e., target sub-tasks); 3) Code Alignment, supporting the Coder and Reviewer to correct the commit by code revision or suggestion feedback to the author; and 4) Document, finalizing by synthesizing the opinions of the CEO, CPO, Coder, and Reviewer to give out the final comments. In addition to six defined roles, the proposed architecture of CodeAgent consists of phase-level and conversation-level components. The waterfall model breaks the code review process at the phase level into four sequential phases. At the conversation level, each phase is divided into atomic conversations. These atomic conversations involve task-oriented role-playing between two agents, promoting collaborative communication. One agent works as an instructor and the other as an assistant. Communication follows an instruction-following style, where agents interact to accomplish a specific subtask within each conversation, and each conversation is supervised by QA-Checker. QA-Checker is used to align the consistency of questions and answers between the instructor and the assistant in a conversation to avoid digression. QA-Checker will be introduced in Section 2.4.
Figure 2 shows an illustrative example of the CodeAgent pipeline. CodeAgent receives the request to do the code review with the submitted commit, commit message, and original files. In the first phase, CEO, CPO, and Coder will cooperate to recognize the modality of input (for example, document, code) and language (for example, Python, Java and Go). In the second phase, with the help of Coder, Reviewer will write an analysis report on consistency analysis, vulnerability analysis, format analysis and suggestions for code revision. Then, in the third phase, according to analysis reports, Coder will align or revise the code if it finds incorrect snippets with the help of Reviewer. cooperates with CTO and CEO to summarize the document and codes about the whole code review in the final phase.
@@ 2.3 Role Card Definition
As shown in Figure 1, we define six characters in our simulation system (CodeAgent), including User, CEO, CPO, CTO, Reviewer, Coder, and they are defined for different specific tasks.
All tasks are processed by the collaborative work of two agents in their multi-round conversations. For example, as a role Reviewer, her responsibility is to do the code review for given codes and files in three aspects (tasks CA, VA, and FA in Sec 2.1) and provide a detailed description of observation. Reviewer’s code review activity is under the assistance with Coder as shown in Figure 2. Meanwhile, with the Reviewer’s assistance, Coder can process the code revision as shown in the ‘Revised codes’ part in the Coder card in Figure 1. Apart from Reviewer, Coder also cooperates with CTO and CEO in the simulated team.
@@ 2.4 Self-Improving CoT with QA Checker
QA-Checker is an instruct-driven agent, designed to fine-tune the question inside a conversation to drive the generated answer related to the question. As shown in Figure 3, the initial question (task instruction) is represented as , and the first answer of the conversation between and is represented as . If QA-Checker identifies that is inappropriate for , it generates additional instructions attached to the original question (task instruction) and combines them to ask agents to further generate a different answer. The combination in Figure 3 is defined as , where is the additional instruction attached. The conversation between two agents is held until the generated answer is judged as appropriate by QA-Checker or it reaches the maximum number of dialogue turns.
@@
The QA-Checker is an instruction-driven agent, crucial in refining questions and answers within a conversation to ensure relevance and precision. Its operation can be understood through the following lemma and proof in Appendix A.
3 Experimental Setup
We evaluate the performance of CodeAgent in various qualitative and quantitative experiments in nine programming languages, on four metrics. In this section, we will discuss experimental settings, including datasets, metrics, and baselines. For more information, please see Appendix C.
@@ 3.1 Datasets
To conduct a fair and reliable comparison for the code revision task, we employ the same datasets (i.e., Trans-Reviewdata, AutoTransformdata, and T5-Reviewdata) as the state-of-the-art study Zhou et al. (2023). Furthermore, we collect and curate an additional dataset targeting the advanced tasks. Table 1 shows our new dataset which includes over 3,545 commits and 2,933 pull requests from more than 180 projects, spanning nine programming languages: Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. It focuses on consistency and format detection, featuring both positive and negative samples segmented by the merged and closed status of pull requests across various languages. The detailed information about the dataset can be seen in Appendix-Section F.
@@ 3.2 Metrics
F1-Score and Recall. We utilized the F1-Score and recall to evaluate our method’s effectiveness on tasks CA and FA. The F1-Score, a balance between precision and recall, is crucial for distinguishing between false positives and negatives. Recall measures the proportion of actual positives correctly identified Hossin and Sulaiman (2015).
Edit Progress (EP). EP evaluates the improvement in code transitioning from erroneous to correct by measuring the reduction in edit distance between the original code and the prediction on task CR. A higher EP indicates better efficiency in code generation Dibia et al. (2022); Elgohary et al. (2021); Zhou et al. (2023).
Hit Rate (Rate) We also use hit rate to evaluate the rate of confirmed vulnerable issues out of the found issues by approaches on task VA.
| CodeBERT | GPT-3.5 | GPT-4.0 | COT | ReAct | CodeAgent | CodeAgent w/o | |
| Find | 1,063 | 864 | 671 | 752 | 693 | 483 | 564 |
| Confirm | 212 | 317 | 345 | 371 | 359 | 449 | 413 |
| Ratecr | 19.94% | 36.69% | 51.42% | 49.34% | 51.80% | 92.96% | 73.23% |
| Rateca | 5.98% | 8.94% | 9.73% | 10.46% | 10.13% | 12.67% | 11.65% |
The values in gray (nn.nn) denote the greatest values for the confirmed number of vulnerabilities and the rates.
@@ 3.3 State-of-the-Art Tools and Models
Our study evaluates various tools and models for code revision and modeling. Trans-Review Tufano et al. (2021) employs src2abs for code abstraction, effectively reducing vocabulary size. AutoTransform Thongtanunam et al. (2022) uses Byte-Pair Encoding for efficient vocabulary management in pre-review code revisions. T5-Review Tufano et al. (2022) leverages the T5 architecture, emphasizing improvement in code review through pre-training on code and text data. In handling both natural and programming languages, CodeBERT Feng et al. (2020) adopts a bimodal approach, while GraphCodeBERT Guo et al. (2021) incorporates code structure into its modeling. CodeT5 Wang et al. (2021), based on the T5 framework, is optimized for identifier type awareness, aiding in generation-based tasks. Additionally, we compare these tools with GPT OPENAI (2022) by OpenAI, notable for its human-like text generation capabilities in natural language processing. Finally, we involve COT Wei et al. (2022) and ReAct Yao et al. (2022), of which COT is a method where language models are guided to solve complex problems by generating and following a series of intermediate reasoning steps and ReAct synergistically enhances language models by interleaving reasoning and action generation, improving task performance and interpretability across various decision-making and language tasks.